# Fast Dynamic 3D Object Generation from a Single-view Video

### v1

@@ Abstract
Generating dynamic three-dimensional (3D) object from a single-view video is challenging due to the lack of 4D labeled data. Existing methods extend text-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling, but they are slow and expensive to scale (e.g., 150 minutes per object) due to the need for back-propagating the information-limited supervision signals through a large pretrained model. To address this limitation, we propose an efficient video-to-4D object generation framework called Efficient4D. It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly train a novel 4D Gaussian splatting model with explicit point cloud geometry, enabling real-time rendering under continuous camera trajectories. Extensive experiments on synthetic and real videos show that Efficient4D offers a remarkable 10-fold increase in speed when compared to prior art alternatives while preserving the same level of innovative view synthesis quality. For example, Efficient4D takes only 14 minutes to model a dynamic object.
11footnotetext:  Corresponding to lizhangfd@fudan.edu.cn
@@ 1 Introduction
Humans possess a remarkable capacity to comprehensively comprehend the spatial and temporal characteristics of a dynamic object in a brief video, even with a limited perspective, enabling them to predict its appearance in unseen viewpoints over time. Despite the significant advancement of three-dimensional (3D) object generation, existing works [28, 17, 4, 41] mostly consider static scenes or objects. With the availability of large-scale 3D datasets [6, 5], training generalizable models capable of directly generating multi-view images becomes possible [20, 21, 22]. These generated images can be turned into a 3D object through reconstruction techniques [25, 39]. By further augmenting these generated static objects with predefined animations [1], dynamic 3D content can be generated. However, this is highly limited due to the need for fine-grained meshes as well as rigid restrictions.
Refer to caption Refer to caption
Refer to caption Refer to caption
Input Generated dynamic objects
@@
Directly generating 4D object/scene content from text description has been recently attempted [36]. To bypass the need of exhaustively labeled training data pairs in form of (text, 4D), it trains a Neural Radiance Fields (NeRF)-like representation [25] via score distillation sampling [28] and separates the modeling of static scene and its dynamics. Not only is this method computationally inefficient caused by heavy supervision back-propagation through a large pretrained model, but also its textual condition is highly ambiguous in expressing the intended visual content. In quest of the aforementioned human’s capability, a recent work [10] proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation. However, similar as [36] this method is also slow to train (e.g., 150 minutes to model a single dynamic object) in addition to complex design, hence unscalable in practice.
To address identified limitations, we formulate an efficient video-to-4D object generation method in a two-staged pipeline called Efficient4D. In the first stage, we generate spacetime-consistent images across different camera views as synthetic training data. This is realized by imposing temporal smoothing filter into a previous multi-view image generator (e.g., SyncDreamer [21]) in tandem with frame interpolation. In the second stage, we use these training data to optimize a novel 4D Gaussian splatting model. This is an extension of the recent 3D Gaussian splatting [12], originally designed for static 3D scene representation, with the temporal dimension introduced additionally on top. Using Gaussian representation brings about further computational efficiency gain along with the explicit geometry property, when compared with NeRF based designs. To tackle the discontinuity between generated frames, we design a inconsistency-aware loss function. A couple of video-to-4D results are presented in Fig. 1.
Our contributions are summarized as follows: (i) We consider for the first time the efficiency challenge with the under-studied video-to-4D object generation problem. (ii) We propose an efficient video-to-4D object generation pipeline, Efficient4D, characterized by directly generating high-quality training data without the need for heavy supervision back-propagation through a large pretrained model as suffered by most 3D/4D object generation approaches. (iii) Instead of using NeRF as the 3D representation model, we introduce a novel 4D Guardian splatting model with the real-time rendering advantage.
(iv) Extensive experiments on both synthetic and real videos validate the significant efficiency advantage (e.g., 10$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the same level of new view synthesis. Additionally, our method can also work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
@@ 2 Related work
Efforts to synthesize videos with free viewpoint control in dynamic scenes have a well-documented history [46]. For example, pre-NeRF [25] approaches faced challenges in reconstructing intricate scene details. Recent advancements in 4D representations, particularly those based on neural rendering, include D-NeRF [29], which decouples geometry and motion, utilizing a canonical space and a learned deformation field. HyperNeRF [27] extends the deformation field to a higher dimension, employing the level-set method for smoother optimization. DeVRF [18] reconstructs the canonical space from multi-view dense observations in the initial frame and learns the deformation field with sparse views in subsequent timesteps, utilizing a voxel grid for representation.
Similarly, deformable 3D Gaussians [44] integrates the deformation field with 3D Gaussian Splatting for the scene’s geometric structure. DyNeRF [14] extends NeRF with an additional latent code to capture detailed temporal variations, while DynIBaR [16] extends the image-based rendering paradigm for representing long videos with complex camera and object motions. Another group of method [7, 3, 33] adopts tensor decomposition of 4D volumes to represent dynamic 3D scenes.
However, these representations often involve complex training schedules or suffer from slow convergence. In this work, drawing inspiration from 3D Gaussian Splatting [12], we propose an efficient 4D scene primitive for the representation of dynamic scenes.
3D generation
3D generation takes two main settigns: text-to-3D and image-to-3D. The pioneering work, DreamFusion [28], introduces the score distillation sampling (SDS) loss for optimizing 3D shapes with diffusion models. SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks [17, 24, 32, 34, 4, 15, 38, 8, 42, 41, 45] and image-to-3D tasks [43, 37, 23, 30] across various dimensions.
However, SDS-based approaches often suffer from extended optimization times. Conversely, some efficient methods [26, 11, 20, 19, 21] have emerged. Notably, Point-E [26] and Shap-E [11] train models to directly generate 3D point clouds or meshes. Zero123 [20] focuses on generating a 2D image from an unseen view based on a single image, convertible to a 3D shape through SDS or [19]. Importantly, SyncDreamer [21] produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
@@
There are a few recent works dedicated for more challenging 4D object generation. For instance, MAV3D [36] deals with a text-to-4D problem by training Hexplane [3] with a video diffusion model and SDS loss. Instead of text input, Consistent4D [10] conditions the generation of 4D object over time on a monocular video with richer and more specific information. However, this work is computationally inefficient due to inheriting the previous SDS loss, along with complex pipeline design. To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the model training speed by 10$\times$ whilst maintaining the quality of novel view synthesis.
Figure 2: Overview of our Efficient4D approach. Given as the input a brief video depicting a dynamic object in limited perspectives, our model aims to generate this object with geometrical and temporal consistency under any specific view and time. Efficient4D comprises two components: (A) Image sequence synthesis across views and timestamps, resulting in (c) an image matrix where each row consists of multi-view geometrically consistent images and each column consists of view-specific temporally consistent images. (B) A novel 4D Gaussian representation model (d) that represents the scene with a number of Gaussian points. It can be trained efficiently and robustly under the confidence-aware (e) supervision on the generated image matrix.
@@ 3 Method
Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video . The model, illustrated in Fig. 2, comprises two key components:
@@ 3.1 Image synthesis across views and timestamps
This synchronization ensures consistent features across $k$ past and $k$ future frames during the denoising process, enhancing temporal consistency.
@@
where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion” as illustrated in Fig. 3. For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency. As $w$ approaches 0.5, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry. This suggests that smoothing at the feature level is effective in aligning frames over time.
However, a challenge arises as the fused features may induce similar motion. While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut. Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence. In practical terms, it is recommended to set the ratio between 0.25 and 0.4, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry.
Refer to caption Refer to caption
Input Feature fusion
@@
Recall that we desire a 3D dynamic object, we should not be content with the discrete images. Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$. For efficient modeling, we formulate a novel 4D Gaussian representation model (Fig. 2(B)), departing from previous slow-to-train 4D reconstruction models [7].
@@
$\mathbf{\mu}_{\mathbf{x}|t}$ $=\mathbf{\mu}_{\mathbf{x}}+\Sigma_{\mathbf{x},t}\Sigma_{t}^{-1}(t-\mathbf{\mu}_{t})$ (7)
@@
During model training, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian. The original objective function, as presented in [12], assumes clean training data, which may not be applicable to synthetic data with inherent imperfections. To address this, we introduce a confidence-aware objective loss formulation defined as:
@@
where $\hat{\mathbf{I}}$ is the unwarped image from adjacent frames estimated by the optical flow method. In such way, the confidence $\mathcal{C}_{\text{rgb/ssim}}$ function as an adaptive role in controlling the loss and gradient by assigning lower weights to inconsistent regions.
@@ 4 Experiments
Figure 4: Qualitative comparisons. We show four images for each method, captured from two viewpoints at two timestamps.
4.1 Implementation details
Competitor
Video-to-4D is a relatively unexplored area, with only one existing method, Consistent4D [10]. In comparison, we obtained the results of Consistent4D by referencing its original paper and project page since the source code has not been publicly released.
Evaluation videos
To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources. Specifically, we focused on 20 sequences: alpaca, astronaut, bird, butterfly, cat, corgi, dinosaur, dog, dragon, duck, flower, guard, panda, penguin, rabbit, robot, snail, Spiderman, yellow face, yoxi.
For dragon, guard, and yoxi, 3D animated models were obtained from Sketchfab [2]. We rendered front-view videos and corresponding masks at 30 fps using Blender. The yellow face sequence was collected from the internet. To facilitate comparison with Consistent4D [10], we also conducted experiments on the videos used in their work [10].
Each sequence consists of monocular videos captured by a stationary camera, with durations ranging from approximately 1s to 6s. We uniformly selected 2-32 frames from each video as keyframes and segmented their foreground parts using SAM [13].
Table 1: Quantitative comparisons. The frame per second (FPS) is evaluated using a $256\times 256$ resolution rendering. ${}^{\dagger}$: Reported in [10] with one V100 GPU. ${}^{\ddagger}$: Measured on K-Planes [7], which partially serves as the fundamental 4D representation of Consistent4D [10].
Method CLIP-similarity $\uparrow$ Generation time $\downarrow$ FPS $\uparrow$ Explicit geometry
Consistent4D [10] 0.8438 165 ${}^{\dagger}$${}^{\ddagger}$
Efficient4D 0.8601 14 min ~1000
Implementation
For time-synchronous spatial volume, we typically set the weight as $\mathbf{w}=(0.1,0.1,0.6,0.1,0.1)$. In the reconstruction stage, we implement a modified CUDA rasterizer tailored to 4D Gaussian. Gaussians are initialized randomly within the space $[-0.6,0.6]^{3}$ with identity rotations and a initial number of 30,000-50,000. Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 9,000 iterations with a batch size of 4. Gaussians are densified every 50 iterations, with densification halted at the 7,000th iteration. All other hyperparameters remain consistent with those in [12]. For speed efficiency, our proposed method only costs about 2 minutes for image generation and 12 minutes for reconstruction on one A6000 GPU.
4.2 Qualitative evaluation
We present qualitative comparisons in Fig. 4. In geometry rendering, we extract points at the mean of Gaussians. We also eliminate white background points and transparent points with low opacity. When assessing texture quality, it is important to note that both methods fall under the category of lifting 2D to 4D. However, Consistent4D produces watercolor-like images with low fidelity, which can be attributed to conflicts from multiple supervisory signals during prolonged optimization. In contrast, our method excels in directly generating high-quality 2D images. Subsequently, our method’s reconstruction stage yields explicit point clouds and exhibits less degradation in rendering quality compared to the SDS-based Consistent4D. Overall, our method consistently outperforms Consistent4D in most cases.

### v2

# https://fudan-zvg.github.io/Efficient4D

@@ Abstract
Generating dynamic 3D object from a single-view video is challenging due to the lack of 4D labeled data. Extending image-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling, existing methods tend to be slow and expensive to scale due to the need for back-propagating the information-limited supervision signals through a large pretrained model. To address this, we propose an efficient video-to-4D object generation framework called Efficient4D. It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly train a novel 4D Gaussian splatting model with explicit point cloud geometry, enabling real-time rendering under continuous camera trajectories. Extensive experiments on synthetic and real videos show that Efficient4D offers a remarkable 20-fold increase in speed when compared to prior art alternatives while preserving the quality of novel view synthesis. For example, Efficient4D takes only 6 mins to model a dynamic object, vs 120 mins by Consistent4D.
Keywords:
4D Generation Gaussian Splatting Video Efficiency
@@ 1 Introduction
Humans possess a remarkable capacity to comprehensively comprehend the spatial and temporal characteristics of a dynamic object in a brief video, even with a limited perspective, enabling them to predict its appearance in unseen viewpoints over time. Despite the significant advancement of 3D object generation, existing works [29, 18, 4, 43] mostly consider static scenes or objects. With the availability of large-scale 3D datasets [6, 5], training generalizable models capable of directly generating multi-view images becomes possible [21, 22, 23]. These generated images can be turned into a 3D object through reconstruction techniques [26, 42]. By further augmenting these generated static objects with predefined animations [1], dynamic 3D content can be generated. However, this approach is still limited due to the need for fine-grained meshes as well as rigid restrictions.
Refer to caption  Refer to caption Refer to caption
 Refer to caption  Refer to caption Refer to caption
Input Consistent4D [13] Efficient4D(Ours)
@@
Directly generating 4D object/scene content from text description has been recently attempted [38]. To bypass the need of exhaustively labeled training data pairs in form of (text, 4D), it trains a Neural Radiance Fields (NeRF)-like representation [26] via score distillation sampling [29] and separates the modeling of static scene and its dynamics. Not only is this method computationally inefficient caused by heavy supervision back-propagation through a large pretrained model, but also its textual condition is highly ambiguous in expressing the intended visual content. In quest of the aforementioned human’s capability, a recent work [13] proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation. However, similar as [38] this method is also slow to train (e.g., 150 minutes to model a single dynamic object) in addition to complex design, hence unscalable and expensive in practice.
To address identified limitations, we formulate an efficient video-to-4D two staged object generation method called Efficient4D. In the first stage, we generate spacetime-consistent images across different camera views as synthetic training data. This is realized by imposing temporal smoothing into a multi-view image generator (e.g., SyncDreamer [22]) in tandem with frame interpolation. In the second stage, we use these training data to optimize a novel 4D Gaussian splatting model. This is an extension of the 3D Gaussian splatting [15], originally designed for static 3D scene representation, with the temporal dimension introduced additionally. Using Gaussian representation brings about further computational efficiency gain along with the explicit geometry property, when compared with NeRF based designs (Fig. 1). To tackle the challenging discontinuity between generated frames, we design an inconsistency-aware loss function.
Our contributions are summarized as follows: (i) We consider for the first time the efficiency challenge with the under-studied video-to-4D object generation problem. (ii) We propose an efficient video-to-4D object generation pipeline, Efficient4D, characterized by directly generating high-quality training data without the need for heavy supervision back-propagation through a large pretrained model as suffered by most 3D/4D object generation approaches. (iii) Instead of using NeRF as the 3D representation model, we introduce a novel 4D Gaussian splatting model with the real-time rendering advantage. (iv) Extensive experiments on both synthetic and real videos validate the significant efficiency advantage (e.g., 20$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the quality of novel view synthesis. Also, our method can work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
@@ 2 Related work
3D generation
3D generation takes two main settigns: text-to-3D and image-to-3D. The pioneering work, DreamFusion [29], introduces the score distillation sampling (SDS) loss for optimizing 3D shapes with diffusion models. SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks [18, 25, 34, 36, 4, 16, 41, 11, 47, 43, 52] and image-to-3D tasks [48, 40, 24, 31] across various dimensions. However, SDS-based approaches often suffer from extended optimization times. Conversely, some efficient methods [27, 14, 21, 20, 22] have emerged. Notably, Point-E [27] and Shap-E [14] train models to directly generate 3D point clouds or meshes. Zero123 [21] focuses on generating a 2D image from an unseen view based on a single image, convertible to a 3D shape through SDS or [20]. Importantly, SyncDreamer [22] produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
@@
Efforts to synthesize videos with free-viewpoint control in dynamic scenes have a well-documented history [53]. For example, pre-NeRF [26] approaches challenges in reconstructing intricate scene details. Recent advancements in 4D representations, particularly those based on neural rendering, include D-NeRF [30], DeVRF [19] and HyperNeRF [28], which decouple geometry and motion, utilizing a canonical space and a learned deformation field. DynIBaR [17] deploys an image-based rendering paradigm for representing long videos with complex camera and object motions. Another group of methods [8, 3, 35] adopt tensor decomposition of 4D volumes to represent dynamic 3D scenes.
Recently, Gaussian Splatting [15] has received widespread attention for its real-time high-fidelity rendering, especially its explicit point-based representation which holds great potential in modeling dynamic scenes. Consequently, a significant amount of work has been proposed to explore its extension to dynamic scene modeling. Among them, Deformable 3D Gaussians [49] and 4DGaussian [46] integrated the deformation field with 3D Gaussian Splatting for the joint learning of the scene’s geometry and dynamics. SC-GS [10] represents the motion with a set of sparse control points to achieve reconstruction and motion editing. Unlike the previous radiance field-based representations often involving complex training schedules or suffering from slow convergence, these methods can be optimized efficiently, achieving real-time rendering while surpassing the past methods in terms of quality.
Our work is perpendicular to all the above works, where any of them can be deployed in our reconstruction phase. But considering both optimization efficiency and expressive capability, we propose to represent dynamic 3D assets by a set of native 4D scene primitives.
@@
There are a few recent works dedicated for more challenging 4D object generation. For instance, MAV3D [38] deals with a text-to-4D problem by training Hexplane [3] with a video diffusion model and SDS loss. Instead of text input, Consistent4D [13] conditions the generation of 4D object over time on a monocular video with richer and more specific information. However, it is computationally inefficient due to inheriting the previous SDS loss, along with complex pipeline design. To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the training speed by 20$\times$ whilst maintaining the quality of novel view synthesis.
@@ 3 Method
Figure 2: Overview of our Efficient4D approach. Given as the input a brief video depicting a dynamic object from a single perspectives, our model aims to generate this object with geometrical and temporal consistency under any specific view and time. Efficient4D comprises two components: (A) Image sequence synthesis across views and timestamps, resulting in (c) an image matrix where each row consists of multi-view geometrically consistent images and each column consists of view-specific temporally consistent images. (B) A novel 4D Gaussian representation model (d) that represents the scene with a number of Gaussian points. It can be trained efficiently and robustly under the confidence-aware (e) supervision on the generated image matrix.
Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video. As illustrated in Fig. 2, it comprises two key components:
@@ 3.1 Image synthesis across views and timestamps
This synchronization ensures consistent features across $k$ past and $k$ future frames during the denoising process, enhancing temporal consistency. With this time-synchronized spatial volumes, the proposed SyncDreamer-T is entirely training-free established on the pretrained SyncDreamer.
@@
Refer to caption Refer to caption
Input Feature fusion
@@
where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion" as illustrated in Fig. 3. For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency. As $w$ approaches 0.5, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry. This suggests that smoothing at the feature level is effective in aligning frames over time.
However, a challenge arises as the fused features may induce similar motion. While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut. Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence. In practical terms, it is recommended to set the ratio between 0.25 and 0.4, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry. Note that the results in Fig. 3 are not sensitive to the choice in this range.
@@
Aiming for 3D dynamic object modeling, discrete images not suffice. Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$. For efficient modeling, we formulate a novel 4D Gaussian representation model (Fig. 2(B)), departing from previous slow-to-train 4D reconstruction models [8].
@@
$\mathbf{\mu}_{\mathbf{x}|t}$ $=\mathbf{\mu}_{\mathbf{x}}+\Sigma_{\mathbf{x},t}\Sigma_{t}^{-1}(t-\mathbf{\mu}_{t}),$ (7)
@@
In training, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian. The original objective function, as presented in [15], assumes clean training data, which may not less valid for synthetic data with inherent imperfections. To address this, we introduce a confidence-aware objective loss formulation defined as:
@@
where $\hat{\mathbf{I}}$ is the unwarped image from adjacent frames estimated by optical flow. In such way, the confidence $\mathcal{C}_{\text{rgb/ssim}}$ function as an adaptive role in controlling the loss and gradient by assigning lower weights to inconsistent regions.
@@ 4 Experiments
4.1 Implementation details
C4D [13]
Ours
GT
Figure 4: Qualitative evaluation of Consistent4D (C4D) [13] and our method against ground truth (GT) on synthetic data.
Competitors
Video-to-4D is a relatively unexplored area, with only one existing method, Consistent4D [13]. In comparison, we obtained the results of Consistent4D by running its released code. We also compare SyncDreamer [22] partially by replacing SyncDreamer-T in Sec. 4.5.
Evaluation videos
To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources. Specifically, we focused on 36 sequences: 32 sequences released by [13] and 4 sequences processed by ourselves. Among the released data, seven sequences are synthetic data where ground truth are provided. Our four sequences are used for sparse input evaluation in Sec. 4.4, which only contain two frames for each sequence. Three of them, named dragon, guard, and yoxi, are rendered from 3D animated models obtained from Sketchfab [2]. The other one named yellow face was collected from the internet. Each sequence consists of monocular videos captured by a stationary camera, with duration ranging from approximately 1s to 3s.
Evaluation metrics
As 4D generation research is still at early stage, there is no well established metric yet. However, we adopt multiple metrics by referring related works for comprehensive evaluation. For evaluating on synthetic data, we use LPIPS score [51] and CLIP similarity [32] between rendered images and ground truth following [13]. For the cases without ground truth, we also use CLIP-similarity between generated images and input frames as [39, 33] to measure image quality. For temporal smoothness, following [50, 7, 9] we use CLIP-T to measure the similarity between adjacent frames of a generated video from different views, including front (CLIP-T-f), side (CLIP-T-s) and back (CLIP-T-b) views. To evaluate a 4D object completely for different methods, we obtain the generated images by rendering 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps.
Implementation
For SyncDreamer-T, we typically set the weight as $\mathbf{w}=(0.1,0.1,0.6,0.1,0.1)$ in time-synchronous spatial volume. The improved sampling strategy [45] can be seamlessly integrated into SyncDreamer-T. In the reconstruction stage, we implement a modified CUDA rasterizer tailored to 4D Gaussian. Gaussians are initialized randomly within the space $[-0.6,0.6]^{3}$ with identity rotations and a initial number of 20,000. Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 20,000 iterations with a batch size of 2. Gaussians are densified every 200 iterations, with densification halted at the 10,000th iteration. All other hyperparameters remain consistent with those in [15]. For speed efficiency, our proposed method only costs about 2 minutes for image generation and 4 minutes for reconstruction on one A6000 GPU.
4.2 Qualitative evaluation
@@
Figure 5: Qualitative comparisons. We show four images per method, captured from two viewpoints at two timestamps. Our geometry is the points at 3D mean of Gaussians.
We first present the results on synthetic data in Fig. 4 with ground truth shown. Our visual results exhibit superior accuracy in both texture and geometry when compared to the ground truth. For other data without ground truth, we present qualitative comparisons in Fig. 5. In geometry rendering, we extract points at the mean of Gaussians. We also eliminate white background points and transparent points with low opacity. When assessing texture quality, it is important to note that both methods fall under the category of lifting 2D to 4D. However, Consistent4D produces watercolor-like images with low fidelity, which can be attributed to conflicts from multiple supervisory signals during prolonged optimization. In contrast, our method excels in directly generating high-quality 2D images. Subsequently, our method’s reconstruction stage yields explicit point clouds and exhibits less degradation in rendering quality compared to the SDS-based Consistent4D. Overall, our method consistently outperforms Consistent4D in most cases.